How does the ideal mechanical advantage (IMA) of an inclined plane vary



with the angle πœƒ of the plane; that is, what trigonometric function of πœƒ gives the value of IMA?

1
Expert's answer
2022-01-14T09:45:47-0500

1) Write the expression for IMA, which is the ration of length of effort over height:


"IMA=\dfrac{L}{H}."


2) Notice that the height can be expressed in terms of length:


"H=L\sin\theta→IMA=\dfrac{L}{L\sin\theta}=\dfrac 1{\sin\theta}=\csc\theta."